OpenStudy (anonymous):

What is the completely factored form of d4 - 8d2 + 16? (d2 + 4)(d2 - 4) (d2 - 4)(d2 - 4) (d2 + 4)(d + 2)(d - 2) (d + 2)(d - 2)(d + 2)(d - 2)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: let z = d^2 so that makes z^2 = d^4 after you square both sides turning d^4 - 8d^2 + 16 into z^2 - 8z + 16

jimthompson5910 (jim_thompson5910):

you now factor z^2 - 8z + 16 to get ________ (fill in the blank)

OpenStudy (anonymous):

(z -4)(z-4) ????

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

now because we made z = d^2, we just replace each copy of z with d^2 to get this final answer (d^2 - 4)(d^2 - 4)

OpenStudy (anonymous):

d^2 or just 2 inside the parethesis?

OpenStudy (anonymous):

or just d*

jimthompson5910 (jim_thompson5910):

It's d^2 or d squared

jimthompson5910 (jim_thompson5910):

not d2 or d times 2

jimthompson5910 (jim_thompson5910):

when you copy/pasted the exponent symbol didn't transfer

OpenStudy (anonymous):

ohhhh i see i see. thankyou!

jimthompson5910 (jim_thompson5910):

np